1. A 9.00 x 102 mL 0.200 M MgI2 was electrolyzed. As a result, hydrogen gas and iodine was formed. 

The volume of hydrogen collected at 260C and 779 mmHg was 1.22 x 103 mL.


(a) Calculate the charge (C) consumed in the process.

(b) How long (in min) did the electrolysis last if a current of 7.55 A was used?


1
Expert's answer
2020-07-17T05:42:38-0400

PV = nRT


So , n = PV / RT .


n =( 779 / 760 ) atm * ( 1.22 * 10-3 ) L . / 0.0821 * ( 299.15 K )


So , number of moles of H2 Gas = 0.0509 mol of H2 Gas .


a ) H2 Gas is Formed By The Reaction 2H+ + 2e ----- H2 ..


So , From the Above Reaction 2 mole Of Electron ( 2 F Or 2 Faraday ) Charge Required for 1 mole of H2 Gas .


So , Charge Required of 0.0509 mol of H2 Gas

=

( 2 * 0.0509 ) F = 2 * 0.0509 * 96500 C = 9,823.7 C . ANSWER .


b ) Charge = Q = Current ( i ) * Time ( t ) .= i * t = ( 7.55 A ) * t . = 9823.7 C .


So , t =( 9823 .7 / 7.55 ) Second = 1,301.152 second . = 21.6 min Answer .


.